# Models of Set Theory: Extensions and Dead-ends

### v5

@@ 6. ADDENDUM AND CORRIGENDUM TO [En-3]
To verify ()($\ast$)( ∗ ), suppose $b_{0}\in B$. By (1), for some $y\in\mathbb{T}$, $\mathrm{V}^{\mathbb{Q}}$ satisfies “$y>_{\mathbb{T}}x_{0}$, $y\in B$, and $f(y)=k$. Therefore there is some $\mathbb{Q}$-condition $q\supseteq q_{0}$ such that:
[…]
Since $B_{0}\in\mathrm{V}$, ()($\ast$)( ∗ ) makes it clear that $B\in\mathrm{V}$ as well, since $B$ can be defined in $\mathrm{V}$ as the downward closure of $B_{0}.$square
  • Let $\Omega$ be the collection consisting of the ranked tree $\tau_{\mathrm{set}}^{\mathcal{N}}$ together with the family of ranked trees $\left\{\tau_{s}:N\models\left|s\right|\geq\aleph_{0}\right\}.$ Since $\mathcal{N}$ is a Rubin model in $\mathrm{V}$, each of the trees in $\Omega$ are rather branchless in $\mathrm{V}$, and Lemma A.7 assures us the ranked trees in $\Omega$ remain rather branchless in $\mathrm{V}^{\mathbb{Q}}$. Lemma A.6, on the other hand, assures us that each branch of the ranked trees in $\Omega$ has a simple definition in $\mathrm{V}^{\mathbb{Q}}.$
  • Let $\mathcal{L}_{\mathrm{set}}^{+}$ be the finite language obtained by augmenting the usual language $\mathcal{L}_{\mathrm{set}}$ of set theory with extra symbols for denoting the binary relation $C$, the binary function $g,$ and the unary function $f$, and let $\mathcal{L}$ be the countable language that is the result of augmenting $\mathcal{L}_{\mathrm{set}}^{+}$ with constants for each element of $\mathcal{M}$. The content of the above bullet item allows us to write a single sentence $\psi$ in $\mathcal{L}_{\mathsf{\omega}_{\mathsf{1}}\mathsf{,\omega}}(\mathsf{Q})$ (where $\mathsf{Q}$ is the quantifier “there exist uncountably many”) that captures the following salient features of the structure $\left(\mathcal{N}^{+},C,g,f\right)$ that hold in $\mathrm{V}^{\mathbb{Q}}$:
@@
    The $\mathcal{L}_{\mathrm{set}}$-reduct $\left(\mathcal{N}^{+},C,g,f\right)$ elementary extends $\mathcal{M}$.
@@
    For each infinite $s$, $\tau_{s}$ has uncountable cofinality, is rather branchless, and $g(s,x)$ weakly specializes $\tau_{s}$.
@@
    $\tau_{\mathrm{set}}^{\mathcal{N}}$ has uncountable cofinality, is rather branchless, and $f$ weakly specializes $\tau_{\mathrm{set}}^{\mathcal{N}}$.
@@
We elaborate on how to write $\psi_{2}$, a similar idea is used to write $\psi_{3}.$ The quantifier Q allows us to express that $\tau_{s}$ has uncountable cofinality, and the range of the specializing function $g(s,x)$ is countable. To expresses the rather branchless feature of $\tau_{s}$, let $\psi(x,y)$ be as in the proof of Lemma A.6. Note that in the notation of Lemma A.6, for any $b\in\mathbb{T}$$\{x\in\mathbb{T}_{s}:\psi(x,b)\}$ is linearly ordered by $\leq_{\mathbb{T}}.$ By Lemma A.6, in $\mathrm{V}^{\mathbb{Q}},$ if $s$ is infinite and $B$ is a branch of $\tau_{s}$, then there is some $b_{0}\in B$ such that $\psi(x,b_{0})$ defines $B$.  Together with the fact that $\tau_{s}$ is rather branchless in $\mathrm{V}^{\mathbb{Q}}$ (thanks to Lemma A.7) the $\mathcal{L}_{\mathsf{\omega}_{\mathsf{1}}\mathsf{,\omega}}$-formula below (no need for $\mathsf{Q}$ here) expresses the rather branchless feature of $\tau_{s}$. In the formula below $\mathbb{T}_{s}$ is the set of the nodes of the ranked tree $\tau_{s}$, and “unbounded rank” is short for “unbounded $\tau_{s}$-rank”.
@@
Step 3 of Stage 2. Let $\psi$ be the sentence described above. Since $\psi$ has a model in $\mathrm{V}^{\mathbb{Q}_{1}\ast\mathbb{Q}}$, and by Keisler’s completeness theorem for $\mathcal{L}_{\omega_{1},\omega}(\mathsf{Q})$ [Ke-2], the consistency of sentences in $\mathcal{L}_{\omega_{1},\omega}(\mathsf{Q})$ is absolute for extensions of $\mathrm{V}$ that do not collapse $\aleph_{1},$ we can conclude that $\psi$ has a model in $\mathrm{V}$ since $\aleph_{1}$ is absolute between $\mathrm{V}$ and $\mathrm{V}^{\mathbb{Q}_{1}\ast\mathbb{Q}}$ (recall that $\mathbb{Q}_{1}$ is $\omega_{1}$-closed and $\mathbb{Q}$ is c.c.c.).)).) . Thanks to Lemmas A.6 and A.7, the $\mathcal{L}_{\mathrm{set}}$-reduct of any model of $\psi$ is a weakly Rubin model that elementarily extends $\mathcal{M}$ (and the cardinality of the model can be arranged to be $\aleph_{1}$ by a Löwenheim-Skolem argument). This concludes the proof of Theorem A.1. $\square$
A.7. Remark. Schmerl [Schm-2] used a similar argument as the one used in the proof of Theorem A.1 to show that every countable model $\mathcal{M}$ of $\mathrm{ZFC}$ has an elementary extension $\mathcal{N}$ of cardinality $\aleph_{1}$ such that every $\omega^{\mathcal{N}}$-complete ultrafilter over $\mathcal{N}$ is definable (and thus coded in $\mathcal{N}$). The proof of Theorem A.1 can be dovetailed with Schmerl’s proof so as to show that every countable model $\mathcal{M}$ of $\mathrm{ZFC}$ has an elementary extension $\mathcal{N}$ of cardinality $\aleph_{1}$ that is weakly Rubin and also has the additional property that every $\omega^{\mathcal{N}}$-complete ultrafilter over $\mathcal{N}$ is definable.

### v6

@@ 1. $(5)$
To verify $(\ast)$, suppose $b_{0}\in B$. By (1), for some $y\in\mathbb{T}$, $\mathrm{V}^{\mathbb{Q}}$ satisfies $y>_{\mathbb{T}}b_{0}$, $y\in B$, and $f(y)=k$. Therefore there is some $\mathbb{Q}$-condition $q\supseteq q_{0}$ such that:
$q\Vdash\left[y>_{\mathbb{T}}x_{0}\wedge y\in B\wedge f(y)=k\right].\vskip 6.0pt plus 2.0pt minus 2.0pt$
Since $B_{0}\in\mathrm{V}$, $(\ast)$ makes it clear that $B\in\mathrm{V}$ as well, since $B$ can be defined in $\mathrm{V}$ as the downward closure of $B_{0}.$ $\square$
  • Let $\Omega$ be the collection consisting of the ranked tree $\tau_{\mathrm{class}}^{\mathcal{N}}$ together with the family of ranked trees $\left\{\tau_{s}^{\mathcal{N}}:\mathcal{N}\models\left|s\right|\geq\aleph_{0}\right\}.$ Since $\mathcal{N}$ is a Rubin model in $\mathrm{V}$, each of the trees in $\Omega$ are rather branchless in $\mathrm{V}$, and Lemma A.7 assures us the ranked trees in $\Omega$ remain rather branchless in $\mathrm{V}^{\mathbb{Q}}$. Lemma A.6, on the other hand, assures us that each branch of the ranked trees in $\Omega$ has a simple definition in $\mathrm{V}^{\mathbb{Q}}.$
  • Let $\mathcal{L}_{\mathrm{set}}^{+}$ be the finite language obtained by augmenting the usual language $\mathcal{L}_{\mathrm{set}}$ of set theory with extra symbols for denoting the binary relation $C$, the binary function $g,$ and the unary function $f$. Then let $\mathcal{L}$ be the countable language that is the result of augmenting $\mathcal{L}_{\mathrm{set}}^{+}$ with constants for each element of $\mathcal{M}$. The content of the above bullet item allows us to write a single sentence $\psi$ in $\mathcal{L}_{\mathsf{\omega}_{\mathsf{1}}\mathsf{,\omega}}(\mathsf{Q})$ (where $\mathsf{Q}$ is the quantifier “there exist uncountably many”) that captures the following salient features of the structure $\left(\mathcal{N}^{+},C,g,f\right)$ that hold in $\mathrm{V}^{\mathbb{Q}}$:
@@ 1. $(1)$
    The $\mathcal{L}_{\mathrm{set}}$-reduct of $\left(\mathcal{N}^{+},C,g,f\right)$ elementarily extends $\mathcal{M}$.
@@ 2. $(2)$
    For each infinite $s$, $\tau_{s}^{\mathcal{N}}$ has uncountable cofinality, is rather branchless, and $g(s,x)$ weakly specializes $\tau_{s}^{\mathcal{N}}$.
@@ 3. $(3)$
    $\tau_{\mathrm{class}}^{\mathcal{N}}$ has uncountable cofinality, is rather branchless, and $f$ weakly specializes $\tau_{\mathrm{class}}^{\mathcal{N}}$.
@@
We elaborate on how to write $\psi_{2}$, a similar idea is used to write $\psi_{3}.$ The quantifier Q allows us to express that $\tau_{s}$ has uncountable cofinality, and the range of the specializing function $g(s,x)$ is countable. To expresses the rather branchless feature of $\tau_{s}$, let $\psi(x,y)$ be as in the proof of Lemma A.6. Note that in the notation of Lemma A.6, for any $b\in\mathbb{T}$ $\{x\in\mathbb{T}_{s}:\psi(x,b)\}$ is linearly ordered by $\leq_{\mathbb{T}}.$ By Lemma A.6, in $\mathrm{V}^{\mathbb{Q}},$ if $s$ is infinite and $B$ is a branch of $\tau_{s}$, then there is some $b_{0}\in B$ such that $\psi(x,b_{0})$ defines $B$. Together with the fact that $\tau_{s}$ is rather branchless in $\mathrm{V}^{\mathbb{Q}}$ (thanks to Lemma A.7) the $\mathcal{L}_{\mathsf{\omega}_{\mathsf{1}}\mathsf{,\omega}}$-formula below (no need for the quantifier $\mathsf{Q}$ here) expresses the rather branchless feature of $\tau_{s}$. In the formula below $\mathbb{T}_{s}$ is the set of the nodes of the ranked tree $\tau_{s}$, and “unbounded rank” is short for “unbounded $\tau_{s}$-rank”.
@@
Step 3 of Stage 2. Let $\psi$ be the sentence described above. Since $\psi$ has a model in $\mathrm{V}^{\mathbb{Q}_{1}\ast\mathbb{Q}}$, and by Keisler’s completeness theorem for $\mathcal{L}_{\omega_{1},\omega}(\mathsf{Q})$ [Ke-2], the consistency of sentences in $\mathcal{L}_{\omega_{1},\omega}(\mathsf{Q})$ is absolute for extensions of $\mathrm{V}$ that do not collapse $\aleph_{1},$ we can conclude that $\psi$ has a model in $\mathrm{V}$ since $\aleph_{1}$ is absolute between $\mathrm{V}$ and $\mathrm{V}^{\mathbb{Q}_{1}\ast\mathbb{Q}}$ (recall that $\mathbb{Q}_{1}$ is $\omega_{1}$-closed and $\mathbb{Q}$ is c.c.c.).). Thanks to Lemmas A.6 and A.7, the $\mathcal{L}_{\mathrm{set}}$-reduct of any model of $\psi$ is a weakly Rubin model that elementarily extends $\mathcal{M}$ (and the cardinality of the model can be arranged to be $\aleph_{1}$ by a Löwenheim-Skolem argument). This concludes the proof of Theorem A.1. $\square$
A.7. Remark. Schmerl [Schm-2] used a similar argument as the one used in the proof of Theorem A.1 to show that every countable model $\mathcal{M}$ of $\mathsf{ZFC}$ has an elementary extension $\mathcal{N}$ of cardinality $\aleph_{1}$ such that every $\omega^{\mathcal{N}}$-complete ultrafilter over $\mathcal{N}$ is $\mathcal{N}$-definable (and thus coded in $\mathcal{N}$). The proof of Theorem A.1 can be dovetailed with Schmerl’s proof so as to show that every countable model $\mathcal{M}$ of $\mathsf{ZFC}$ has an elementary extension $\mathcal{N}$ of cardinality $\aleph_{1}$ that is weakly Rubin and also has the additional property that every $\omega^{\mathcal{N}}$-complete ultrafilter over $\mathcal{N}$ is $\mathcal{N}$-definable.